OpenStudy (anonymous):

log3 (27)/ log3 (9)=27/9...Is the statement true or false and if so how?

OpenStudy (anonymous):

seems unlikely

OpenStudy (anonymous):

I've came up with two different answers 3 and 0.

OpenStudy (anonymous):

\(\log_3(27)=3\) since \(3^3=27\) also \(\log_3(9)=2\) since \(3^2=9\) therefore \[\frac{\log_3(27)}{\log_3(9)}=\frac{3}{2}\]

OpenStudy (anonymous):

wow ok thank you